Calculate the period of a microwave that has a speed of 3 x 10^8 m/s and a wavelength of 300 x 10^-2 m


1
Expert's answer
2021-04-23T11:10:57-0400
"T=\frac{\lambda}{v}\\T=\frac{3}{3\cdot10^8}=10^{-8}\ s"